OpenStudy (anonymous):

Using the formulae (provided below) to find the derivative of y=arcsec(x/a), where a is constant.

OpenStudy (anonymous):

\[d/dx arcsec(u)=\frac{ 1 }{ u \sqrt{u^{2}-1} } du/dx\]